OpenStudy (anonymous):

What is the 46th term of the arithmetic series 1, 8, 15, … ?

OpenStudy (amistre64):

a series is different from a sequence. a series is the sum of a sequence.

OpenStudy (amistre64):

are you being asked to find the 46th partial sum? or are you being asked to find the 46th term of this sequence?

OpenStudy (amistre64):

S1 = a1 = 1 S2 = a1 + a2 = 8 S3 = a1 + a2 + a3 = 15 a1 = 1 a2 = 7 a3 = 7 does not exhibit an arithmatic sequence

OpenStudy (amistre64):

so most likely you are askng for the 46th term of an arithmatic sequence